OpenStudy (anonymous):

2 sin^2 theta= sin theta. Find 4 solutions please!

OpenStudy (cwrw238):

first write it as 2 sin^2 theta - sin theta = 0 now factor left hand side

OpenStudy (anonymous):

Not sure how this would factor out

OpenStudy (cwrw238):

sin theta is common to both terms so take it out and put the rest in parentheses

OpenStudy (anonymous):

so it would be 2sin theta(sin theta- sin theta)?

OpenStudy (cwrw238):

no its sin theta(2 sin theta - 1 ) = 0

OpenStudy (cwrw238):

you divide 2 sin theta - sin theta by sin theta and the result goes into the parentheses

OpenStudy (cwrw238):

now when the product of 2 functions = 0 either one of them must = 0 so you have to solve the following 2 equations sin theta = 0 and 2 sin theta - 1 = 0

OpenStudy (anonymous):

I get how to solve it now but I am just a little confused on how you came up with the 2 equations from that one equation

OpenStudy (cwrw238):

consider a simpler example x(x - 2) = 0 x times (x - 2) = 0 for the result to be zero either x = 0 or (x - 2) = 0

OpenStudy (cwrw238):

any number times zero equals zero

OpenStudy (anonymous):

Spot on. Perfect explanation! thank you for the help!

OpenStudy (cwrw238):

yw